# CollabStory: Multi-LLM Collaborative Story Generation and Authorship Analysis

### v1

@@ Abstract
The rise of unifying frameworks that enable seamless interoperability of Large Language Models (LLMs) has made LLM-LLM collaboration for open-ended tasks a possibility. Despite this, there have not been efforts to explore such collaborative writing. We take the next step beyond human-LLM collaboration to explore this multi-LLM scenario by generating the first exclusively LLM-generated collaborative stories dataset called CollabStory. We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author stories. We generate over 32k stories using open-source instruction-tuned LLMs. Further, we take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for human-human multi-author writing tasks and analysis. We extend their authorship-related tasks for multi-LLM settings and present baselines for LLM-LLM collaboration. We find that current baselines are not able to handle this emerging scenario. Thus, CollabStory is a resource that could help propel an understanding as well as the development of techniques to discern the use of multiple LLMs. This is crucial to study in the context of writing tasks since LLM-LLM collaboration could potentially overwhelm ongoing challenges related to plagiarism detection, credit assignment, maintaining academic integrity in educational settings, and addressing copyright infringement concerns. We make our dataset and code available at https://github.com/saranya-venkatraman/multi_llm_story_writing.
@@ 1 Introduction
Dataset # Stories # Authors  Avg words M-M Collaboration Available
STORIUM Akoury et al. (2020) 5,743 30,119 ~19k tokens H-H X
CoAuthor Lee et al. (2022) 830 58 418 H-M X
StoryWars Du and Chilton (2023) 40,135 9,494 367 H-H X X
CollabStory [Ours] 32,503 5 725 M-M
[…]
@@
We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author creative stories. This exploration is novel, as previous studies have primarily focused on human-LLM collaboration. Towards this goal, we generate the first multi-LLM collaborative story dataset called CollabStory using open-source LLMs. We select 5 frequently used LLMs (with number of downloads on HuggingFace for May 2024 provided in parenthesis): Meta’s Llama ($>540k$ downloads, Touvron et al. (2023)), Mistral.ai’s Mistral ($>1000k$ downloads, Jiang et al. (2023)), Google’s Gemma ($>180k$ downloads, Team et al. (2024)), AllenAI’s Olmo ($>26k$ downloads, Groeneveld et al. (2024)) and Microsoft’s Orca ($>22k$ downloads, Mukherjee et al. (2023)) to replicate a scenario in which commonly used LLMs from different organizations are being used in conjunction towards a single task. We demonstrate how one such dataset can be developed and the considerations involved in building an iterative Multi-LLM story-writer. We take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for multi-author writing tasks and analysis for human-human collaboration for over 15 years. We replicate their task settings and present baselines of different authorship-related tasks such as authorship verification and attribution for LLM-LLM collaboration and demonstrate that current baselines are challenged by this emerging scenario. CollabStory is the first resource that could help propel an understanding as well as the development of new techniques to discern the use of multiple LLMs in text.
Our work is motivated by the implications of Multi-LLM settings for different stakeholders (LLM developers, end-users) and considerations (such as credit assignment, legality of usage) arising in the generative AI landscape. As one example, a malicious actor might assemble texts from different LLMs together in one document to evade current detectors and successfully spread misinformation. Our discussion will further elaborate on the tasks our dataset enables and why it is crucial to develop methods to tackle the incoming challenges of machine-machine collaboration.
@@ 2 Related Work
Despite such emerging tools, only a handful have developed datasets that can be leveraged to understand collaborative story writing. One such resource is the STORIUM dataset Akoury et al. (2020) that contains over 5k creative stories written and obtained from human-human collaboration. In terms of human-machine co-writing, CoAuthor Lee et al. (2022) and CoPoet Chakrabarty et al. (2022) remain one of the few publically available datasets of human-machine collaborative creative story and poem writing, respectively. Beyond creative writing, Zeng et al. (2024) developed the first machine-human academic essay dataset as a means to study boundary detection for academic settings. A comparison of CollabStory with existing datasets is provided in Table 1.
@@
Table 2: Summary of Data Collection Statistics
@@ 3.1 CollabStory: Dataset Creation
We generate a dataset of creative stories using five open-source instruction-tuned LLMs: Llama2 Touvron et al. (2023), Olmo Groeneveld et al. (2024), Gemma Team et al. (2024), Mistral Jiang et al. (2023) and Orca Mukherjee et al. (2023) (model details are provided in Table 2). The main focus of our data generation is to simulate a scenario where LLMs from different sources (organizations) collaboratively work on a storyline, handing off control of the story from one LLM to the next. The stories in our dataset vary in the number of authors/LLMs involved, from being written entirely by a single LLM to written collaboratively by between 2 to up to all 5 LLMs. In this document, we refer to each of the LLMs as “authors". For cases where we refer to the human author, we specifically mention “human” author/writer. We generate our dataset by prompting various LLMs using creative writing prompts from an existing dataset called the Writing Prompts (WP) Dataset. The Writing Prompts Dataset was collected by Fan et al. (2018) using Reddit’s r/WritingPrompts/ forum that contains premises or prompts for stories. The WP dataset consists of a cleaned subset of story prompts and corresponding human-written stories using filtration criteria such as removal of stories that are bot-generated, less than 30 words long, contain profanity, general announcements, and so on. We used the test split [^4] https://www.kaggle.com/code/ratthachat/writingprompts-combine-one-line-data-for-gpt2/input?select=writingPrompts of this dataset as the source of prompts for LLM generated stories. We also filter out prompts that do not have at least one corresponding human-written story that is at least 800 words long. We do this to ensure that the prompt itself does not preclude longer storylines. We chose 800 words as a criteria as a means to include stories that are slightly longer than the average of the dataset. The average length (number of words) of articles in the test set is 675.75 words. Out of 15138 total prompt-story pairs, this left us with 4623 data points. For each prompt, we divide the total goal article length (800-900 words) by the number of authors ($N$) to calculate the length of each part or story chunk to be written by each author, such that the writing load is distributed roughly uniformly amongst the LLM authors. We also generate different permutations of LLM authorship order such that every author can contribute to random parts of the story and we ensure that our dataset does not have any spurious correlations between LLM/author and story sections such as the beginning, or ending. For each value of the numbers of authors i.e. $N\in\{1,2,3,4,5\}$, we generate all possible permutations of author orders. For example, for  N = 3, two examples of author order permutations could be:
@@ 3.2 LLM prompting
For each value of $N$, we used different prompts to generate story parts sequentially, as detailed in Table 3. Utilizing the vLLM library [^5] https://docs.vllm.ai/en/stable/, we accessed and generated text from various LLMs. Initially, we conducted a pilot study to refine our prompts by generating and reviewing 100 articles. For the "Beginning" prompt, the first LLM used only the original r/WritingPrompts/ input. For subsequent parts, we found that longer input prompts reduced story length, so we used Falcon.ai summarizer [^6] https://huggingface.co/Falconsai to condense the story so far into under 80 words, allowing LLMs to generate longer sequences. We also included the last sentence of the story so far for smooth continuity. Prompts for different sections only varied in their instructions to “begin”, “continue”, or “conclude” the story. Additionally, we added an instruction to prevent LLMs from generating extraneous instructions. More details are provided in Section A.2.
Story Part Prompt Template
Beginning You are a creative story writer. Write a story that starts with the prompt {starting prompt} in around {n} words. Do not add any instructions. Start the story as follows:
Middle Write {n} words to continue this storyline: {summary of story so far}. Continue from this sentence: {last sentence from previous part}
Ending Write {n} words to conclude this storyline: {summary of story so far}. Do not add any instructions. Continue from this sentence: {last sentence from previous part}
@@ 3.3 Post-processing and filtering
Feature N=1 (H) N=1 (M) N=2 N=3 N=4 N=5
# Words 1352.26 ± 425.11 725.03 ± 288.32 1090.67 ± 207.43 1154.44 ± 112.24 1091.99 ± 85.67 995.42 ± 74.20
# Sentences 84.23 ± 34.85 41.90 ± 61.94 60.49 ± 14.49 64.38 ± 11.57 59.29 ± 11.24 53.86 ± 9.41
# Words in sentence 17.82 ± 28.18 18.27 ± 3.92 18.42 ± 2.85 18.30 ± 2.45 18.81 ± 2.47 18.88 ± 2.51
Vocabulary richness 0.34 ± 0.05 0.40 ± 0.09 0.36 ± $0.05^{*}$ 0.36 ± $0.03^{*}$ 0.37 ± $0.03^{*}$ 0.39 ± 0.03
% of stopwords 31.26 ± 4.76 37.00 ± 4.77 37.77 ± 3.11 37.39 ± 2.67 37.71 ± 2.39 37.41 ± 2.39
Readability 80.28 ± 33.26 75.16 ± 9.83 75.34 ± 8.13 75.25 ± 7.80 74.37 ± 8.00 74.80 ± 8.14
Entropy 38.49 ± 12.25 26.03 ± 13.10 38.42 ± 7.93 40.30 ± 4.75 37.97 ± 4.07 34.69 ± 3.58
Coherence 0.38 ± 0.05 0.46 ± 0.06 0.46 ± 0.06 0.46 ± 0.04 0.46 ± 0.04 0.46 ± 0.04
Table 5: Comparison of descriptive features for articles with number of authors $N\in[1,5]$, as compared with human-written single-author stories for the same prompts. Here, “H” $\to$ “Human”; “M”$\to$ Machine. * represents statistical significance ($\textit{p}<0.01$) compared to human-written stories.
3.4 Descriptive Statistics
@@
We report the average and standard deviation of number of words, sentences, vocabulary richness, readability scores, and coherence scores using TextDescriptives Library [^7] https://github.com/HLasse/TextDescriptives for all parts of the story, as well as for different numbers of authors ($N$) in Table 5. There is a statistically significant difference between LLM-coauthored and human-written stories in vocabulary richness measured by type-token-ratio (TTR) for $N\in[2,3,4]$. For all other measures, our dataset’s machine-generated stories follow similar distributions to human-written stories. We thus ensure the quality of our dataset using automated measures of readability and coherence by using human-generated stories as the reference text. Detailed statistics for each LLM are provided in the Appendix (Table 12).
@@ 4.1 Task 1: Is a story written by multiple authors or not?
We randomly sample articles from the single-LLM authored stories i.e. $N=1$ as the negative class v/s articles from the multi-authored settings where $N\in[2,3,4,5]$ as the positive class. We sample from the single-LLM stories to keep the class distribution equal, based on the number of articles for each $N$. From Table 6, we see that for all methods, the performance at $N=5$ is higher than for $N=2$, gradually increasing with the value of $N$. Stories that have a higher number of authors are more distinct from single-authored ones. We conjecture that introducing more authors in the article might lead to more variations in the text, making stories with $N=5$ authors most easily distinguishable from stories without any such variations i.e. $N=1$.
Task 1: Multi-author or not
Method N=2 N=3 N=4 N=5
MNB 0.83 0.86 0.86 0.87
SVM 0.78 0.8 0.83 0.82
BERT 0.85 0.85 0.83 0.87
ALBERT 0.83 0.89 0.9 0.92
RoBERTa 0.88 0.92 0.9 0.93
AVG 0.83 0.86 0.86 0.88
Table 6: Performance is shown as F1-scores. AVG denotes average F1-score for each $N$. For each $N\geq 2$ we evaluate the classifiers on their ability to distinguish the stories from those written with $N=1$. Best performing method is in bold and second highest underlined for each N.

### v2

@@ Abstract
The rise of unifying frameworks that enable seamless interoperability of Large Language Models (LLMs) has made LLM-LLM collaboration for open-ended tasks a possibility. Despite this, there have not been efforts to explore such collaborative writing. We take the next step beyond human-LLM collaboration to explore this multi-LLM scenario by generating the first exclusively LLM-generated collaborative stories dataset called CollabStory. We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author stories. We generate over 32k stories using open-source instruction-tuned LLMs. Further, we take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for human-human multi-author writing tasks and analysis. We extend their authorship-related tasks for multi-LLM settings and present baselines for LLM-LLM collaboration. We find that current baselines are not able to handle this emerging scenario. Thus, CollabStory is a resource that could help propel an understanding as well as the development of techniques to discern the use of multiple LLMs. This is crucial to study in the context of writing tasks since LLM-LLM collaboration could potentially overwhelm ongoing challenges related to plagiarism detection, credit assignment, maintaining academic integrity in educational settings, and addressing copyright infringement concerns. We make our dataset and code available at https://anonymous.4open.science/r/multi_llm_story_writing-55AE/README.md.
@@ 1 Introduction
Dataset # Stories # Authors Avg # Words M-M Collaboration Available
STORIUM Akoury et al. (2020) 5,743 30,119 ~19k tokens H-H
CoAuthor Lee et al. (2022) 830 58 418 H-M
StoryWars Du and Chilton (2023) 40,135 9,494 367 H-H
CollabStory [Ours] 32,503 5 725 M-M
Table 1: Comparison of CollabStory with other existing collaborative creative story datasets. Here, “M-M” $\to$“Machine-Machine” ,“H-H” $\to$“Human-Human”, where “H” $\to$ “Human” and “M” $\to$ Machine. Ours is the largest dataset to present multi-LLM or machine-machine collaborative generation.
@@
We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author creative stories. This exploration is novel, as previous studies have primarily focused on human-LLM collaboration. Towards this goal, we generate the first multi-LLM collaborative story dataset called CollabStory using open-source LLMs. We select 5 frequently used LLMs (with number of downloads on HuggingFace for May 2024 provided in parenthesis): Meta’s Llama ($>540k$ downloads, Touvron et al. (2023)), Mistral.ai’s Mistral ($>1000k$ downloads, Jiang et al. (2023)), Google’s Gemma ($>180k$ downloads, Team et al. (2024)), AllenAI’s Olmo ($>26k$ downloads, Groeneveld et al. (2024)) and Microsoft’s Orca ($>22k$ downloads, Mukherjee et al. (2023)) to replicate a scenario in which commonly used LLMs from different organizations are being used in conjunction towards a single task. We demonstrate how one such dataset can be developed and the considerations involved in building an iterative Multi-LLM story-writer. We present an evaluation of the generated texts and their quality as the number of authors increases from one to five. Since this is the first such multi-LLM dataset, our evaluation provides insights into the current state of LLMs and their ability to collaborate on open-ended generation tasks. We then take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for multi-author writing tasks and analysis for human-human collaboration for over 15 years. We adapt their authorship-related task suite to the machine-machine settings for problems such as authorship verification and attribution, and demonstrate that current baselines are challenged by this emerging scenario. CollabStory is the first resource that could help propel an understanding as well as the development of new techniques to discern the use of multiple LLMs in text.
Our work is motivated by the implications of Multi-LLM settings for different stakeholders (LLM developers, end-users) and considerations (such as credit assignment, legality of usage) arising in the generative AI landscape. As an example, a malicious actor might assemble texts from different LLMs together in one document to evade attribution and successfully spread misinformation. Our findings elaborate on the tasks our dataset enables and emphasize the importance of tackling the incoming challenges of machine-machine collaboration.
@@ 2 Related Work
Despite such emerging tools, only a handful have developed datasets that can be leveraged to understand collaborative story writing. One such resource is the STORIUM dataset Akoury et al. (2020) that contains over 5k creative stories written and obtained from human-human collaboration. In the case of human-machine co-writing, CoAuthor Lee et al. (2022) and CoPoet Chakrabarty et al. (2022) remain one of the few publically available datasets of human-machine collaborative creative story and poem writing, respectively. Beyond creative writing, Zeng et al. (2024) developed the first machine-human academic essay dataset as a means to study boundary detection for academic settings. A comparison of CollabStory with existing datasets is provided in Table 1.
@@
Table 2: CollabStory creation: Summary of generation statistics
@@ 3.1 CollabStory: Dataset Creation
We generate a dataset of creative stories using five open-source instruction-tuned LLMs: Llama2 Touvron et al. (2023), Olmo Groeneveld et al. (2024), Gemma Team et al. (2024), Mistral Jiang et al. (2023) and Orca Mukherjee et al. (2023) (model details are provided in Table 2). The main focus of our data generation is to simulate a scenario where LLMs from different sources (organizations) collaboratively work on a storyline, handing off control of the story from one LLM to the next. The stories in our dataset vary in the number of authors/LLMs involved, from being written entirely by a single LLM to written collaboratively by between 2 to up to all 5 LLMs. In this document, we refer to each of the LLMs as “authors". For cases where we refer to the human author, we specifically mention “human” author/writer. We generate our dataset by prompting various LLMs using creative writing prompts from an existing dataset called the Writing Prompts (WP) Dataset. The Writing Prompts Dataset was collected by Fan et al. (2018) using Reddit’s r/WritingPrompts/ forum that contains premises or prompts for stories. The WP dataset consists of a cleaned subset of story prompts and corresponding human-written stories using filtration criteria such as removal of stories that are bot-generated, less than 30 words long, contain profanity, general announcements, and so on. We used the test split [^4] https://www.kaggle.com/code/ratthachat/writingprompts-combine-one-line-data-for-gpt2/input?select=writingPrompts of this dataset as the source of prompts for LLM generated stories.
We also filter out prompts that do not have at least one corresponding human-written story that is at least 800 words long. We do this to ensure that the prompt itself does not preclude longer storylines. We chose 800 words as a criteria as a means to include stories that are slightly longer than the average of the dataset. The average length (number of words) of articles in the test set is 675.75 words. Out of 15138 total prompt-story pairs, this left us with 4623 data points. For each prompt, we divide the total goal article length (800-900 words) by the number of authors ($N$) to calculate the length of each part or story chunk to be written by each author, such that the writing load is distributed roughly uniformly amongst the LLM authors. We also generate different permutations of LLM authorship order such that every author can contribute to random parts of the story and we ensure that our dataset does not have any spurious correlations between LLM/author and story sections such as the beginning, or ending. For each value of the numbers of authors i.e. $N\in\{1,2,3,4,5\}$, we generate all possible permutations of author orders. For example, for  N = 3, two examples of author order permutations could be:
@@ 3.2 LLM prompting
For each value of $N$, we used different prompts to generate story parts sequentially, as detailed in Table 3. Utilizing the vLLM library [^5] https://docs.vllm.ai/en/stable/, we accessed and generated text from various LLMs. Initially, we conducted a pilot study to refine our prompts by generating and reviewing 100 articles. For the "Beginning" prompt, the first LLM used only the original r/WritingPrompts/ input. For subsequent parts, we found that longer input prompts reduced story length, so we used the Falcon.ai summarizer [^6] https://huggingface.co/Falconsai to condense the story so far into under 80 words, allowing LLMs to generate longer sequences. We also included the last sentence of the story so far for smooth continuity. Prompts for different sections only varied in their instructions to begin, continue, or conclude the story. Additionally, we added an instruction to prevent LLMs from generating extraneous instructions. More details are provided in Section A.1.
Prompt Templates
Beginning Prompt
You are a creative story writer. Write a story that starts with the prompt {starting prompt} in around {n} words. Do not add any instructions. Start the story as follows:
Middle Prompt
Write {n} words to continue this storyline: {summary of story so far}. Continue from this sentence: {last sentence from previous part}
Ending Prompt
Write {n} words to conclude this storyline: {summary of story so far}. Do not add any instructions. Continue from this sentence: {last sentence from previous part}
@@ 3.3 Post-processing and filtering
Figure 2: N on the X-axis denotes the number of authors, and N=1(H) and N=1(M) correspond to the human-written and machine-generated single-authored texts, respectively. All other texts (N >=2) are multi-LLM generated. Y-axis shows the values of the measure shown in each suplot as mentioned in the headings. For all measures, we show average and standard deviation for N going from 1 to 5. For all measures except vocabulary richness (3rd column, 1st row), increasing the number of authors (N) does not lead to statistically significant deviations from the human text distribution.
Figure 3: F1 scores for authorship-related tasks in the multi-LLM scenario, using different methods (color-coded) as the number of authors increases from N=1 to N=5.
4 Dataset Analysis
@@
We compare the LLM-generated single and multi-authored texts with the single-authored human written stories to study the relative quality of LLM generated stories as the number of authors present increases. To do this, we report the average and standard deviation of different measures such as the average number of (i) words, (ii) sentences, (iii) vocabulary richness, (iv) percentage of stopwords, (v) readability scores, (vi) entropy, and (vii) coherence scores using the TextDescriptives Library [^7] https://github.com/HLasse/TextDescriptives for different numbers of authors ($N$) in Figure 2. We measure (vii) creativity using OCSAI (Open Creativity Scoring with Artificial Intelligence) Organisciak et al. (2023) that provides text creativity scores using an LLM (GPT-4) fine-tuned on human annotations of creativity. From Figure 2, we see that across most measures, there isn’t a significant deviation or decline as the number of authors is increased, i.e. collaboration or introducing multiple LLMs in the generation setting does not disturb the quality of the stories generated. In fact, the only statistically significant difference between LLM-coauthored and human-written stories in vocabulary richness measured by type-token-ratio (TTR) for $N\in[2,3,4]$. For all other measures, our dataset’s machine-generated stories follow similar distributions to human-written stories. Automated measures of readability and coherence utilized the human-generated stories as the reference text, and even for these measures, we see that collaboratively generated text scores do not deviate from the reference score distribution. This analysis indicates that LLMs are able to collaboratively generate stories without any significant changes in quality through sequential prompting. Detailed statistics for each LLM are provided in the Appendix (Table 7).
Figure 4: To evaluate the continuity of story parts, we sample two potential continuations to a story part, one correct and one incorrect. In this figure, we see an example of a story part on the left and two candidates to be evaluated for continuity. In green are some terms highlighted to show continuity in topics, characters, and plot lines, and in red are highlighted sections of the text that suggest discontinuity.
Negative Sample Source Correct Wins Neither Wins Incorrect Wins
Different Story 75% 25% 0%
Same Story 35% 55.33% 9.67%
Table 5: Continutity evaluation of pairs of story parts using GPT-4o using different negative sample sources (either from within the same story or not). This table presents the percentage of story parts for which either only the correct part was evaluated as having continuity, or the negative sample (incorrect) or both.
4.1 Story Continuity
An important aspect of collaborative storywriting is the notion of continuity. Particularly, we wanted to measure if the different story parts generated sequentially by different LLM authors followed a logical and cohesive plot. To evaluate this, we used a prompt-based evaluation using GPT-4o OpenAI (2024) to discern if two consecutive story parts continue a story well or not. To do this, we develop two types of pairs of story parts: one pair containing the "Correct" or the positive sample which points to the actual next part in the story, or "Incorrect" which is the negative sample that is either a random story part drawn from a different story or from the same one (but not the true next one, any other part). In other words, for the negative samples taken from within the same story, we ensure to not use the correct consecutive part as a negative sample. An example of this setup is illustrated in Figure 4, where given a story part, there is an actual continuation ("Correct") and a random negative sample ("Incorrect") story part drawn from either the same story or a different one. We then use the following prompt structure to obtain continuity evaluations for both candidate next parts and then compare if the actual continuation was the one with a better continuation evaluation. We provide GPT-4o a pair of texts at a time (out of a total of 600 pairs) and ask it the following question: “Does Part 2 serve as a good continuation of Part 1 in terms of logical flow, coherence, and consistency? Please respond with Yes or No.” GPT-4o generated the binary "yes/no" responses and an explanation. We then compare the percentage of story parts for which either the correct answer got a "Yes" and the incorrect one got a "No" response (Correct Wins), if both candidate continuations were evaluated the same (Neither Wins), or if the Incorrect answer got a "Yes" and the Correct one got a "No". We present the results in Table 5. We see that when the negative story parts are sampled from different stories than the one that the first part belongs to, then there are 0 cases for which the incorrect part wins. This is an expected result, as story parts that belong to different stories would have different topics, plot lines, characters, and so on. Hence, detecting this discontinuity should not be hard, as is the case. An example of a case where Neither Wins and both candidate story parts get an evaluation as being discontinuous is provided in Figure 4. As illustrated by this example, it is crucial to note that our prompt encourages the model to evaluate the logical or common-sense flow of the stories. Hence, topic overlap alone was insufficient to mark two parts as continuous. In the harder setting, when the negative sample is also a story part from within the same story, we see that there is a marked increase in both the Neither Wins and Incorrect Wins scenarios (see Table 5) where Incorrect Wins goes from 0 to 9.67% and Neither Wins has an over 30% increase. This is expected since many of the story parts were not the next immediate part but a few sequences apart. They all follow the same story and share some logical flow despite the distance in their occurrences in the story. For example, the fourth part of story might get a positive continuity evaluation, just as the second, third, and fifth would if compared with the first part of the story. Over 85% of Neither Wins story pairs had a continuous evaluation for both candidates (142 out of 166). Thus, out-of-the-box LLMs are able to follow plot lines and logic even when continuing each others’ partially written parts at any point in the story.
@@ 5.1 Task 1: Is a story written by multiple authors or not?
We randomly sample articles from the single-LLM authored stories i.e. $N=1$ as the negative class v/s articles from the multi-authored settings where $N\in[2,3,4,5]$ as the positive class. We sample from the single-LLM stories to keep the class distribution equal, based on the number of articles for each $N$. From Figure 3(a), we see that for all methods, the performance at $N=5$ is higher than for $N=2$, gradually increasing with the value of $N$. Stories that have a higher number of authors are more distinct from single-authored ones. We conjecture that introducing more authors in the article might lead to more variations in the text, making stories with $N=5$ authors most easily distinguishable from stories without any such variations i.e. $N=1$.